# Computation of Smyth and Hoare Power Constructions in Well-filtered Dcpos

### v2

@@ 5 The double power well-filtered dcpo
(1): To prove that $L$ is a $\mathcal{Q}$-algebra, let $\alpha^{\mathcal{Q}}=\alpha\circ\mathcal{Q}\eta_{L}:\mathcal{Q}(L)\rightarrow L$. By Lemma 5.10, we know that $\alpha^{\mathcal{Q}}$ will be the corresponding structure map as long as $\mathcal{Q}\eta$ is a morphism of monads from $\mathcal{Q}$ to $\mathcal{QH}$. As it happens, it is Theorem 5.3 together with Proposition 5.4 that guarantees this fact. So $(L,\alpha^{\mathcal{Q}})$ is a $\mathcal{Q}$-algebra.
@@
We proceed by defining $\alpha^{\mathcal{H}}=\alpha\circ\theta_{\mathcal{H}(L)}:\mathcal{H}(L)\rightarrow L$. Utilizing Theorem 5.3 and Proposition 5.4 again, we can get that $\theta_{\mathcal{H}}$ is a morphism of monads from $\mathcal{H}$ to $\mathcal{QH}$. So we conclude that $(L,\alpha^{\mathcal{H}})$ is an $\mathcal{H}$-algebra.
(2): By the fact that for each object $P$ in $\mathbf{WF}$, $(\mathcal{Q}(P),\iota_{P})$ is a $\mathcal{Q}$-algebra, we know $(\mathcal{QH}(L),\iota_{\mathcal{H}(L)})$ is also a $\mathcal{Q}$-algebra. Then according to the definition of $\mathcal{Q}$-algebra homomorphism, we need to show the diagram
$\mathcal{Q}^{2}\mathcal{H}(L)$$\mathcal{QH}(L)$$\mathcal{Q}(L)$$L$$\iota_{\mathcal{H}(L)}$$\alpha^{\mathcal{Q}}$$\mathcal{Q}\alpha$$\alpha$
commutes, that is, $\alpha^{\mathcal{Q}}\circ\mathcal{Q}\alpha=\alpha\circ\iota_{\mathcal{H}(L)}$. Note that
$\alpha^{\mathcal{Q}}\circ\mathcal{Q}\alpha$ $=\alpha\circ\mathcal{Q}\eta_{L}\circ\mathcal{Q}\alpha$
$=\alpha\circ\mathcal{Q}(\eta_{L}\circ\alpha)$
$=\alpha\circ\mathcal{Q}(\mathcal{H}\alpha\circ\eta_{\mathcal{QH}(L)})\quad\mathrm{by\ naturality\ of\ \eta}$
$=\alpha\circ\mathcal{QH}\alpha\circ\mathcal{Q}\eta_{\mathcal{QH}(L)}$
$=\alpha\circ\rho_{L}\circ\mathcal{Q}\eta_{\mathcal{QH}(L)}\qquad\mathrm{since\ (L,\alpha)\ is\ a\ \mathcal{QH}{-}algebra},$
which means that we just need to show
$\rho_{L}\circ\mathcal{Q}\eta_{\mathcal{QH}}=\iota_{\mathcal{H}(L)}:\mathcal{Q}^{2}\mathcal{H}(L)\rightarrow\mathcal{QH}(L)$.
Assume $\mathbf{K}\in\mathcal{Q}^{2}\mathcal{H}(L)$. One can calculate that $\iota_{\mathcal{H}(L)}(\mathbf{K})=\bigcup\mathbf{K}$ and
$\rho_{L}\circ\mathcal{Q}\eta_{\mathcal{QH}}(\mathbf{K})={\uparrow}\{\bigcup\mathcal{C}:\mathcal{C}\in\bigcup{\uparrow}\{\phi_{\mathcal{H}}(\mathbf{A}):\mathbf{A}\in{\uparrow}\{{\downarrow}\{\mathcal{K}\}:\mathcal{K}\in\mathbf{K}\}\}\}$,
in which we take $\rho=\mathcal{Q}\mu\circ\rho_{\mathcal{HH}}\circ\mathcal{Q}\phi_{\mathcal{H}}$. On the one hand, for any $C\in\bigcup\mathbf{K}$, $C\in\mathcal{K}_{0}$ for some $\mathcal{K}_{0}\in\mathbf{K}$, which means ${\downarrow}\{C\}\bigcap\mathcal{K}_{0}\neq\emptyset$, thus for each $\mathcal{M}\supseteq\mathcal{K}_{0}$, i.e., $\mathcal{M}\in{\downarrow}\{\mathcal{K}_{0}\}$, we have ${\downarrow}\{C\}\bigcap\mathcal{M}\neq\emptyset$. It follows that ${\downarrow}\{C\}\in\phi_{\mathcal{H}}({\downarrow}\{\mathcal{K}_{0}\})$, and further, ${\downarrow}\{C\}\in\bigcup{\uparrow}\{\phi_{\mathcal{H}}(\mathbf{A}):\mathbf{A}\in{\uparrow}\{{\downarrow}\{\mathcal{K}\}:\mathcal{K}\in\mathbf{K}\}\}$. Then the fact that $C=\bigcup{\downarrow}\{C\}$ implies $C\in\rho_{L}\circ\mathcal{Q}\eta_{\mathcal{QH}}(\mathbf{K})$. So $\iota_{\mathcal{H}(L)}(\mathbf{K})\subseteq\rho_{L}\circ\mathcal{Q}\eta_{\mathcal{QH}}(\mathbf{K})$ holds. On the other hand, for each $\mathcal{C}\in\bigcup{\uparrow}\{\phi_{\mathcal{H}}(\mathbf{A}):\mathbf{A}\in{\uparrow}\{{\downarrow}\{\mathcal{K}\}:\mathcal{K}\in\mathbf{K}\}\}$, there exist an $\mathbf{M}\in\mathcal{QHH}(L)$ and an $\mathbf{A}_{0}\in{\uparrow}\{{\downarrow}\{\mathcal{K}\}:\mathcal{K}\in\mathbf{K}\}$ such that $\mathcal{C}\in\mathbf{M}\subseteq\phi_{\mathcal{H}}(\mathbf{A}_{0})$. Note that $\mathbf{A}_{0}\in{\uparrow}\{{\downarrow}\{\mathcal{K}\}:\mathcal{K}\in\mathbf{K}\}$ means that $\mathcal{K}_{0}\in\mathbf{A}_{0}$ for some $\mathcal{K}_{0}\in\mathbf{K}$, thus $\mathcal{C}\bigcap\mathcal{K}_{0}\neq\emptyset$. Suppose that $C_{0}$ is in this intersection. Then $C_{0}\subseteq\bigcup\mathcal{C}$; moreover, by the saturation of $\mathcal{K}_{0}$, $\bigcup\mathcal{C}\in\mathcal{K}_{0}$, which implicates $\bigcup\mathcal{C}\in\bigcup\mathbf{K}$. Hence, $\rho_{L}\circ\mathcal{Q}\eta_{\mathcal{QH}}(\mathbf{K})\subseteq\iota_{\mathcal{H}(L)}(\mathbf{K})$.
As argued above, we have $\rho_{L}\circ\mathcal{Q}\eta_{\mathcal{QH}}(\mathbf{K})=\iota_{\mathcal{H}(L)}(\mathbf{K})$ and conclude that $\alpha$ is a $\mathcal{Q}$-algebra homomorphism.
As for $\alpha\circ\phi$, since $(L,\alpha^{\mathcal{H}})$ and $(\mathcal{HQ}(L),\mu_{\mathcal{Q}(L)})$ are both $\mathcal{H}$-algebras and by the definition of $\mathcal{H}$-algebra homomorphism, we need to verify that the diagram below commutes,
$\mathcal{H}^{2}\mathcal{Q}(L)$$\mathcal{HQ}(L)$$\mathcal{H}(L)$$L$$\mu_{\mathcal{Q}(L)}$$\alpha^{\mathcal{H}}$$\mathcal{H}(\alpha\circ\phi)$$\alpha\circ\phi$
more precisely, $\alpha^{\mathcal{H}}\circ\mathcal{H}(\alpha\circ\phi)=\alpha\circ\phi\circ\mu_{\mathcal{Q}(L)}$ holds. One can see that
$\alpha^{\mathcal{H}}\circ\mathcal{H}(\alpha\circ\phi)$ $=\alpha\circ\theta_{\mathcal{H}(L)}\circ\mathcal{H}\alpha\circ\mathcal{H}\phi$
$=\alpha\circ\mathcal{QH}\alpha\circ\theta_{\mathcal{HQH(L)}}\circ\mathcal{H}\phi\quad\mathrm{by\ naturality\ of\ \theta}$
$=\alpha\circ\rho_{L}\circ\theta_{\mathcal{HQH(L)}}\circ\mathcal{H}\phi.$
Thus it is enough to show $\rho_{L}\circ\theta_{\mathcal{HQH(L)}}\circ\mathcal{H}\phi=\phi\circ\mu_{\mathcal{Q}(L)}$. The verification process is a routine similar to that of $\rho_{L}\circ\mathcal{Q}\eta_{\mathcal{QH}}=\iota_{\mathcal{H}(L)}$. Hence, we could draw the conclusion that $\alpha\circ\phi$ is an $\mathcal{H}$-algebra homomorphism.
[…]
$k({\uparrow}\{A\})=k\circ\theta_{\mathcal{H}(L)}(A)=m\circ\theta_{\mathcal{H}(L)}(A)=m({\uparrow}\{A\})$,
which induces
$k(\mathcal{K})=k(\bigcup_{A\in\mathcal{K}}{\uparrow}\{A\})=\bigcup_{A\in\mathcal{K}}k({\uparrow}\{A\})=\bigcup_{A\in\mathcal{K}}m({\uparrow}\{A\})=m(\bigcup_{A\in\mathcal{K}}{\uparrow}\{A\})=m(\mathcal{K})$
for each $\mathcal{K}\in\mathcal{QH}(L)$, that is to say, $k=m$. Note that $\alpha\circ\theta_{\mathcal{H}(L)}=\alpha^{\mathcal{H}}=\beta^{\mathcal{H}}=\beta\circ\theta_{\mathcal{H}(L)}$. Then the uniqueness of $m$ enables us to conclude that $\alpha=\beta$.
@@
If the distributive law $\phi$ is an isomorphism and $L$ is a $\mathcal{QH}$-algebra, then $L$ is a frame.
@@
The fact given in Proposition 5.11 that $L$ is a $\mathcal{Q}$-algebra and an $\mathcal{H}$-algebra yields that $L$ is indeed a meet continuous complete lattice by Corollary 3.8 and Proposition 3.16. Thus the proof will be completed if we show that $L$ is distributive.
@@
It follows from Proposition 5.11 that $\alpha$ is a $\mathcal{Q}$-homomorphism. Thus $\alpha$ preserves finite infs by Lemma 3.15. For any $\mathcal{K}_{1},\mathcal{K}_{2}\in\mathcal{QH}(L)$, since $\phi$ is surjective, there exist $\mathcal{A}_{1},\mathcal{A}_{2}\in\mathcal{HQ}(L)$ such that $\phi(\mathcal{A}_{i})=\mathcal{K}_{i},i=1,2$. Thus
$\alpha(\mathcal{K}_{1}\vee\mathcal{K}_{2})$ $=\alpha(\mathcal{K}_{1}\cap\mathcal{K}_{2})=\alpha(\phi(\mathcal{A}_{1})\cap\phi(\mathcal{A}_{2}))=\alpha(\phi(\mathcal{A}_{1}\cup\mathcal{A}_{2}))$
$=\alpha\circ\phi(\mathcal{A}_{1}\vee\mathcal{A}_{2})=\alpha\circ\phi(\mathcal{A}_{1})\vee\alpha\circ\phi(\mathcal{A}_{2})$
$=\alpha(\mathcal{K}_{1})\vee\alpha(\mathcal{K}_{2}),$
where the penultimate equation holds thanks to the fact that $\alpha\circ\phi$ is an $\mathcal{H}$-homomorphism proved in Proposition 5.11.
@@
One can calculate that $\alpha\circ\varphi_{1}\circ(\theta_{\mathcal{H}}\circ\eta_{L})^{3}(x,y,z)=x\wedge(y\vee z)$ and $\alpha\circ\varphi_{2}\circ(\theta_{\mathcal{H}}\circ\eta_{L})^{3}(x,y,z)=(x\wedge y)\vee(x\wedge z)$. It is obvious that $\varphi_{1}=\varphi_{2}$ by Claim 1. Thus $x\wedge(y\vee z)=(x\wedge y)\vee(x\wedge z)$, that is, $L$ satisfies the distributive law. In conclusion, $L$ is a frame. ∎
@@
Let $f:(L,\alpha)\rightarrow(M,\beta)$ be a $\mathcal{QH}$-algebra homomorphism. It was shown in Theorem 5.3 that $\mathcal{Q}\eta:\mathcal{Q}\rightarrow\mathcal{QH}$ and $\theta_{\mathcal{H}}:\mathcal{H}\rightarrow\mathcal{QH}$ are morphisms of monads; hence, by Lemma 5.10, we know $f:(L,\alpha\circ\mathcal{Q}\eta_{L})\rightarrow(M,\beta\circ\mathcal{Q}\eta_{M})$ is a $\mathcal{Q}$-algebra homomorphism and $f:(L,\alpha\circ\theta_{\mathcal{H}(L)})\rightarrow(M,\beta\circ\theta_{\mathcal{H}(L)})$ is an $\mathcal{H}$-algebra homomorphism. It follows that $f$ preserves finite infs and arbitrary sups, that is to say, $f$ is a frame homomorphism.
@@
If the distributive law $\phi$ is an isomorphism, then the Eilenberg-Moore category of $\mathcal{QH}$ over $\mathbf{WF}$ is a subcategory of the category $\mathbf{Frm}$.

### v3

@@ 5 The double power well-filtered dcpo
By Proposition 2.4 (1) and (2), $(\mathcal{Q}\mathcal{H}(L),\rho_{L})$ is a $\mathcal{Q}\mathcal{H}$-algebra and $\alpha:(\mathcal{Q}\mathcal{H}(L),\rho_{L})\rightarrow(L,\alpha)$ is a $\mathcal{Q}\mathcal{H}$-algebra homomorphism.
  1. 1.
    Theorem 5.3 and Proposition 5.5 imply that $\mathcal{Q}\eta:\mathcal{Q}\rightarrow\mathcal{Q}\mathcal{H}$ is a morphism of monads. By Lemma 5.4, $(\mathcal{Q}\mathcal{H}(L),\rho_{L}\circ\mathcal{Q}\eta_{L})$ and $(L,\alpha\circ\mathcal{Q}\eta_{L})=(L,\alpha^{\mathcal{Q}})$ are $\mathcal{Q}$-algebras, and $\alpha:(\mathcal{Q}\mathcal{H}(L),\rho_{L}\circ\mathcal{Q}\eta_{L})\rightarrow(L,\alpha^{\mathcal{Q}})$ is a $\mathcal{Q}$-algebra homomorphism. By Proposition 2.4 (1), $(\mathcal{Q}\mathcal{H}(L),\iota_{\mathcal{H}(L)})$ is also a $\mathcal{Q}$-algebra. Since $\mathcal{Q}$ is a right $\rm{KZ}$-monad, Proposition 5.8 indicates that structure maps for $\rm{KZ}$-monads are unique, $\rho_{L}\circ\mathcal{Q}\eta_{L}=\iota_{\mathcal{H}(L)}$ holds; therefore, $\alpha:(\mathcal{Q}\mathcal{H}(L),\iota_{\mathcal{H}(L)})\rightarrow(L,\alpha^{\mathcal{Q}})$ is a $\mathcal{Q}$-algebra homomorphism.
  2. 2.
@@
    Theorem 5.3 and Proposition 5.5 imply that $\theta_{\mathcal{H}}:\mathcal{H}\rightarrow\mathcal{Q}\mathcal{H}$ is a morphism of monads. By Lemma 5.4, $(\mathcal{Q}\mathcal{H}(L),\rho_{L}\circ\theta_{\mathcal{H}(L)})$ and $(L,\alpha\circ\theta_{\mathcal{H}(L)})=(L,\alpha^{\mathcal{H}})$ are $\mathcal{H}$-algebras, and $\alpha:(\mathcal{Q}\mathcal{H}(L),\rho_{L}\circ\theta_{\mathcal{H}(L)})\rightarrow(L,\alpha^{\mathcal{Q}})$ is an $\mathcal{H}$-algebra homomorphism.
  3. 3.
    Assume that $\beta:\mathcal{QH}(L)\rightarrow L$ is also a structure map of $L$. Then $\beta^{\mathcal{Q}}=\beta\circ\mathcal{Q}\eta_{L}$ and $\beta^{\mathcal{H}}=\beta\circ\theta_{\mathcal{H}(L)}$ are $\mathcal{Q}$-algebra and $\mathcal{H}$-algebra structure maps, respectively. Since both $\mathcal{Q}$ and $\mathcal{H}$ are $\rm{KZ}$-monads, it follows from Proposition 5.8 that each $\mathcal{Q}$-algebra ($\mathcal{H}$-algebra) has at most one structure map. Thus $\beta^{\mathcal{Q}}=\alpha^{\mathcal{Q}}$ and $\beta^{\mathcal{H}}=\alpha^{\mathcal{H}}$. By (1), $\alpha$ and $\beta$ are $\mathcal{Q}$-algebra homomorphisms from $(\mathcal{Q}\mathcal{H}(L),\iota_{\mathcal{H}(L)})$ to $(L,\alpha^{\mathcal{Q}})=(L,\beta^{\mathcal{Q}})$. Since $\alpha\circ\theta_{\mathcal{H}(L)}=\alpha^{\mathcal{H}}=\beta^{\mathcal{H}}=\beta\circ\theta_{\mathcal{H}(L)}$, Proposition 2.4 (3) implies $\alpha=\beta$.
@@
If $L$ is a $\mathcal{QH}$-algebra, then $L$ is a frame.
@@
The fact given in Proposition 5.11 that $L$ is a $\mathcal{Q}$-algebra and an $\mathcal{H}$-algebra yields that $L$ is indeed a meet continuous complete lattice by Corollary 3.8 and Proposition 3.16. Thus the proof will be completed if we prove that $L$ is distributive.
@@
By Proposition 5.11, $\alpha$ is both a $\mathcal{Q}$-homomorphism and an $\mathcal{H}$-homomorphism. Thus $\alpha$ preserves finite infs by Lemma 3.15 and all sups by Lemma 3.7.
@@
One can calculate that $\alpha\circ\varphi_{1}\circ(\theta_{\mathcal{H}}\circ\eta_{L})^{3}(x,y,z)=x\wedge(y\vee z)$ and $\alpha\circ\varphi_{2}\circ(\theta_{\mathcal{H}}\circ\eta_{L})^{3}(x,y,z)=(x\wedge y)\vee(x\wedge z)$. It is obvious that $\varphi_{1}=\varphi_{2}$ by Claim 1. Thus $x\wedge(y\vee z)=(x\wedge y)\vee(x\wedge z)$, that is, $L$ satisfies the distributive law. In conclusion, $L$ is a frame.
@@
Let $f:(L,\alpha)\rightarrow(M,\beta)$ be a $\mathcal{QH}$-algebra homomorphism. It was shown in Theorem 5.3 that $\mathcal{Q}\eta:\mathcal{Q}\rightarrow\mathcal{QH}$ and $\theta_{\mathcal{H}}:\mathcal{H}\rightarrow\mathcal{QH}$ are morphisms of monads; hence, by Lemma 5.4, we know $f:(L,\alpha\circ\mathcal{Q}\eta_{L})\rightarrow(M,\beta\circ\mathcal{Q}\eta_{M})$ is a $\mathcal{Q}$-algebra homomorphism and $f:(L,\alpha\circ\theta_{\mathcal{H}(L)})\rightarrow(M,\beta\circ\theta_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{H}(M)}})$ is an $\mathcal{H}$-algebra homomorphism. It follows that $f$ preserves finite infs and arbitrary sups, that is to say, $f$ is a frame homomorphism.
@@
The Eilenberg-Moore category of $\mathcal{QH}$ over $\mathbf{WF}$ is a subcategory of the category $\mathbf{Frm}$.
@@
Remark 5.16.
The general observation that the algebras of the double power construction are frames is due to S. Vickers [24] or perhaps P. Taylor [23], although Vickers works with locales and Taylor works in an abstract category of locally compact sober spaces. In an abstract category of locally compact sober spaces, Taylor proved that when the power constructions commute, the composite monad is equivalent to the monad defined as double exponentiation by the Sierpinski space, and characterized its algebras as frames. Furthermore, the characterizations work in more general categories when the necessary exponentials exist. We know every core-compact well-filtered dcpo is locally compact sober [17] and by Corollary 4.9, for each locally compact sober dcpo, the Hoare and the Smyth power constructions commute with each other. Thus, our composite monad $\mathcal{Q}\mathcal{H}$ is also equivalent to the monad defined as double exponentiation by the Sierpinski space in the core-compact case. Then a natural open question is whether this equivalence extends to a more general category of well-filtered dcpos satisfying conditions (i) and (ii) of Theorem 4.9. We leave this for future investigation.